OpenStudy (anonymous):

What is 4x+2y=20 in slope-intercept form ?

OpenStudy (anonymous):

the equation for slope intercept form: y = mx + b 4x + 2y = 20 2y = 20 -4x dividing both side by 2 y = 10 - 2x therefore slope intercept form is y = -2x + 10 answer.

OpenStudy (anonymous):

y=-2x+10